HFSC(8) Linux HFSC(8) NAME HFSC−HierarchicalFairServiceCurve’scontrolunderlinux SYNOPSIS tcqdiscadd...hfsc[defaultCLASSID] tcclassadd...hfsc[[rtSC][lsSC]|[scSC]][ulSC] rt:realtimeservicecurve ls:linkshareservicecurve sc:rt+lsservicecurve ul:upperlimitservicecurve •atleastoneofrt,lsorscmustbespecified •ulcanonlybespecifiedwithlsorsc SC:=[[m1BPS]dSEC]m2BPS m1:slopeofthefirstsegment d:x−coordinateofintersection m2:slopeofthesecondsegment SC:=[[umaxBYTE]dmaxSEC]rateBPS umax:maximumunitofwork dmax:maximumdelay rate:rate FordescriptionofBYTE,BPSandSEC−pleaseseeUNITSsectionoftc(8). DESCRIPTION(qdisc) HFSCqdischasonlyoneoptionalparameter−default.CLASSIDspecifiestheminorpartof thedefaultclassid,wherepacketsnotclassifiedbyothermeans(e.g.u32filter,CLASSIFYtarget ofiptables)willbeenqueued.Ifdefaultisnotspecified,unclassifiedpacketswillbedropped. DESCRIPTION(class) HFSCclassisusedtocreateaclasshierarchyforHFSCscheduler.Forexplanationofthealgorithm,andthemeaningbehindrt,ls,scandulservicecurves−pleaserefertotc−hfsc(7). AsyoucanseeinSYNOPSIS,servicecurve(SC)canbespecifiedintwoways.Eitherasmaximumdelayforcertainamountofwork,orasabandwidthassignedforcertainamountoftime. Obviously,m1issimplyumax/dmax. Bothm2andratearemandatory.Ifyouomitotherparameters,youwillspecifylinearservice curve. SEEALSO tc(8),tc−hfsc(7),tc−stab(8) Pleasedirectbugreportsandpatchesto:<net...@vger.kernel.org> AUTHOR ManpagecreatedbyMichalSoltys(sol...@ziu.info) iproute2 31October2011 1